OpenStudy (anonymous):

What is the simplest form of the expression? 3 (sq rt)750 + 3 (sq rt)2058 - 3 (sq rt)48 Please help

OpenStudy (anonymous):

a. 14 (sq rt) 6 b. 17.5 (sq rt) 6 c. 33 (sq rt) 6 d. 10 (sq rt) 6

OpenStudy (anonymous):

\[3\sqrt{750}+3\sqrt{2058}-3\sqrt{48}\] The difficult part is finding perfect square factors of these large numbers. The first one is easy because 25 goes in evenly. The second one is much more difficult. I used a calculator and tried perfect squares: 4, then 9, I can see that 16 won't go in if 4 doesn't and I can see that 25 does not go in, so the next one is 36, and then 49 - that one worked The last one was quicker as 16 went in evenly. So this is what I have now: \[3\sqrt{25*30}+3\sqrt{49*42}-3\sqrt{16*3}\]

OpenStudy (anonymous):

Thank you, but that doesn't really help my question.?

OpenStudy (anonymous):

we can pull apart multiplication (or division but NOT addition or subtraction) \[3\sqrt{25*30}=3\sqrt{25}*\sqrt{30}=3*5\sqrt{30}=15\sqrt{30}\]

OpenStudy (anonymous):

You have to do the same thing with the other 2 terms

OpenStudy (anonymous):

I dont see how any of those answers will work

OpenStudy (anonymous):

I don't see how you got your answer.. I'm very confused.

OpenStudy (anonymous):

Is this the original: \[3\sqrt{750}+3\sqrt{2058}+3\sqrt{48}\]

OpenStudy (anonymous):

I know what the original is..

OpenStudy (anonymous):

I am checking because I am wondering if you have a copy error - ultimately I get: \[15\sqrt{30}+21\sqrt{42}+12\sqrt{3}\] You cannot simplify further and that was not one of your choices so I am thinking there is a copy error somewhere

OpenStudy (anonymous):

I think those are supposed to be indexes